# Data Attribution for Text-to-Image Models by Unlearning Synthesized Images

### v1

@@ 5.1 Leave-$K$-out counterfactual evaluation
Figure 6: Customized Model benchmark [18]. We report Recall@10 (left) and mAP (right) and show performance on artist-style models (y-axis) vs. object-centric models (x-axis). On object-style models, our method performs on par with AbC features, which were directly tuned on the benchmark, while significantly outperforming them on artist-style models. We plot 1 standard error on both axes.
Figure 7: Qualitative examples on the Customized Model benchmark. The red boxes indicate ground truth exemplar images used for customizing the model. Both our method and baselines successfully identify the exemplar images on object-centric models (left), while our method outperforms the baselines with artistic style models (right).
5.2 Customized Model Benchmark
Wang et al. [18] focus on a specialized form of attribution: attributing customized models trained on an individual or a few exemplar images. This approach provides ground truth attribution since the images generated by customized models are computationally influenced by exemplar images. While this evaluation has limited generalization to attribution performance with larger training sets, it is the only tractable evaluation for attributing large-scale text-to-image models to date.
Evaluation protocol.   Since the Customized Model Benchmark has ground truth, the problem is evaluated as a retrieval task. We report Recall@K and mAP, measuring the success of retrieving the exemplar images amongst a set including 100K LAION images. We compare with Wang et al.’s feature-matching approach that finetunes on the Customized Model dataset, referred to as DINO (AbC) and CLIP (AbC). For our evaluation, we selected a subset of the dataset comprising 20 models: 10 object-centric and 10 artist-style models. We select 20 synthesized images with different prompts for each model, resulting in 400 synthesized image queries.
Comparing with AbC features.   We report Recall@10 and mAP in Figure 6. Our method performs on par with baselines when testing on object-centric models, while significantly outperforming them on artist-style models. Although CLIP (AbC) and DINO (AbC) are fine-tuned for this attribution task, the feature matching approach can sometimes confuse whether to attribute a synthesized image to style-related or object-related images. In contrast, our method, which has access to the model itself, traces influential training images more effectively. In Figure 7, we show a qualitative example. While DINO (AbC) and CLIP (AbC) can retrieve visually or semantically similar images, our method successfully identifies the exemplars in both cases. We include ablation studies in Appendix C.2.
@@ 6 Discussion, Broader Impacts, and Limitations
Our work proposes a method for data attribution for text-to-image models, leveraging model unlearning. We provide a counterfactual validation, verifying that removing the identified influential images indeed destroys the target image. While our method empirically demonstrates that unlearning can be effectively used, work remains to make this practical. Though our model unlearns efficiently, estimating the reconstruction loss on the training set remains a bottleneck, as several forward passes are required on each training estimate. While our evaluation showed that unlearning is useful for attribution, direct evaluation of unlearning algorithms for large generative models remains an open research challenge. Furthermore, to find a critical set of images, our method and baselines assign influence scores to individual images and sort them. However, groups of images may have interactions that are not captured in such a system. Furthermore, our method and baselines explore attribution of the whole image, while finer attribution on individual aspects of the image, such as style, structure, or individual segments, are of further interest.
@@ A.3 Connection to influence functions
$\mathcal{L}(\mathbf{z},\theta_{-\hat{\mathbf{z}}})$ $\approx\mathcal{L}(\mathbf{z},\theta_{0})+\nabla\mathcal{L}(\mathbf{z},\theta_{0})(\theta_{-\hat{\mathbf{z}}}-\theta_{0})$ (15)
$=\mathcal{L}(\mathbf{z},\theta_{0})+\gamma\nabla\mathcal{L}(\mathbf{z},\theta_{0})F^{-1}\nabla\mathcal{L}(\hat{\mathbf{z}},\theta).$
@@
$\approx\gamma\nabla\mathcal{L}(\mathbf{z},\theta_{0})F^{-1}\nabla\mathcal{L}(\hat{\mathbf{z}},\theta).$
In this special case, our method is equivalent to influence function $\nabla\mathcal{L}(\mathbf{z},\theta_{0})F^{-1}\nabla\mathcal{L}(\hat{\mathbf{z}},\theta)$, after approximations. Practically, the difference between our method and influence functions is that we are taking larger, sometimes multiple steps (rather than a single, infinitesimal step), and are explicitly evaluating the loss (rather than with a linear, closed-form approximation).
@@ C.1 MSCOCO Models
Deviation of generated output in leave-$K$-out models.   We report the quantitative evaluation for the deviation of generated output in terms of MSE in Figure 8 and in terms of CLIP similarity in Figure 9. We find that in terms of this metric, our method still outperforms all baselines while being slightly less performant than $K=4000$ when compared with DINO features under CLIP similarity. Interestingly, we find that JourneyTRAK performs better than TRAK in this metric despite significantly underperforming regarding loss changes (Figure 3). We note that DINO, despite not being designed for attribution, clearly outperforms both TRAK and JourneyTRAK in this evaluation setup.
@@
We also compare different subsets of weights to optimize and report loss change, deviation measured by MSE, and deviation measured by CLIP similarity in Figure 121314, respectively. We find that the 4 choices of weight subset selection all lead to effective attribution performance, where restricting weight updates to cross-attention KV yields the best performance overall. However, both configurations for SGD updates perform much worse, indicating the importance of regulating unlearning with Fisher information.
@@ C.2 Customized Model Benchmark
Refer to caption Refer to caption

### v2

@@ 5.2 Customized Model Benchmark
Figure 6: Customized Model benchmark [18]. We report Recall@10 (left) and mAP (right) and show performance on artist-style models (y-axis) vs. object-centric models (x-axis). On object-centric models, our method performs on par with AbC features, which were directly tuned on the benchmark, while significantly outperforming them on artist-style models. D-TRAK performs the second best on artist-style models but worse on object-centric models. We plot one standard error on both axes.
@@ 6 Discussion, Broader Impacts, and Limitations
Our work proposes a method for data attribution for text-to-image models, leveraging model unlearning. We provide a counterfactual validation, verifying that removing the identified influential images indeed destroys the target image. While our method empirically demonstrates that unlearning can be effectively used, work remains to make this practical. Though our model unlearns efficiently, estimating the reconstruction loss on the training set remains a bottleneck, as several forward passes are required on each training estimate, as profiled in Appendix D. While our evaluation showed that unlearning is useful for attribution, direct evaluation of unlearning algorithms for large generative models remains an open research challenge. Furthermore, to find a critical set of images, our method and baselines assign influence scores to individual images and sort them. However, groups of images may have interactions that are not captured in such a system. Furthermore, our method and baselines explore attribution of the whole image, while finer attribution on individual aspects of the image, such as style, structure, or individual segments, are of further interest.
@@ A.3 Connection to influence functions
$\mathcal{L}(\mathbf{z},\theta_{-\hat{\mathbf{z}}})$ $\approx\mathcal{L}(\mathbf{z},\theta_{0})+\nabla\mathcal{L}(\mathbf{z},\theta_{0})^{T}(\theta_{-\hat{\mathbf{z}}}-\theta_{0})$ (15)
$=\mathcal{L}(\mathbf{z},\theta_{0})+\gamma\nabla\mathcal{L}(\mathbf{z},\theta_{0})^{T}F^{-1}\nabla\mathcal{L}(\hat{\mathbf{z}},\theta).$
@@
$\approx\gamma\nabla\mathcal{L}(\mathbf{z},\theta_{0})^{T}F^{-1}\nabla\mathcal{L}(\hat{\mathbf{z}},\theta).$
In this special case, our method is equivalent to influence function $\nabla\mathcal{L}(\mathbf{z},\theta_{0})^{T}F^{-1}\nabla\mathcal{L}(\hat{\mathbf{z}},\theta)$, after approximations. Practically, the difference between our method and influence functions is that we are taking larger, sometimes multiple steps (rather than a single, infinitesimal step), and are explicitly evaluating the loss (rather than with a linear, closed-form approximation).
@@ B.2 Customized Model Benchmark
D-TRAK   We follow the best-performing hyperparameter reported in D-TRAK, using a random projection dimension of 32768 and lambda of 500. We use a single model to compute the influence score.
@@ C.1 MSCOCO Models
Additional plots.   We visualize the baseline comparisons in Table 1 as plots, where Figure 12,13,14 represents loss change, MSE deviation, and CLIP deviation, respectively. We also visualize ablations studies in Table 2 as plots (Figure 15,16,17).
[…]
@@
We also compare different subsets of weights to optimize and report plots for loss change, deviation measured by MSE, and deviation measured by CLIP similarity in Figure 151617, respectively.
Target Images (to forget) Other Images (to retain)
MSE ($\uparrow$) CLIP ($\downarrow$) MSE ($\downarrow$) CLIP ($\uparrow$)
SGD 0.081±0.003 0.67±0.01 0.033±0.0004 0.83±0.002
Full weight 0.086±0.005 0.70±0.01 0.039±0.0010 0.86±0.002
Ours 0.093±0.004 0.65±0.01 0.022±0.0004 0.89±0.002
Table 3: Effectiveness in Unlearning Synthesized Images. We compare different choices of unlearning algorithms and evaluate based on whether the method can forget the target images and retain other images. We measure the performance with regenerated images’ deviations via mean square error (MSE) and CLIP similarity. SGD refers to the naive baseline without EWC regularization, and full weight refers to updating on all of the weights instead of cross-attention KV.
Figure 7: Ablation for unlearning. We find that our unlearning method (Ours) outperforms other variants (Full Weight, SGD) in terms of forgetting the target (left) while retaining other concepts (right).
Effectiveness in unlearning synthesized images.   Our attribution method relies on unlearning synthesized images, making it crucial to have an unlearning algorithm that effectively removes these images without forgetting other concepts. We analyze the performance of our unlearning algorithm itself and ablate our design choices. We construct experiments by unlearning a target synthesized image and evaluating:
  • Unlearning the target image: We measure the deviation of the regenerated image from the original model’s output—the greater the deviation, the better.
  • Retaining other concepts: We generate 99 images using different text prompts and evaluate their deviations from the original model’s output—the smaller the deviation, the better.
We measure these deviations using mean square error (MSE) and CLIP similarity. We evaluate across 40 target images, with text prompts sampled from the MSCOCO validation set. We compare to the following ablations:
  • SGD refers to swapping our method’s Newton update steps (Eq. 5 ) to the naive baseline, where we run SGD steps to maximize the target loss without EWC regularization.
  • Full weight refers to running Newton update steps on all of the weights instead of Cross Attn KV.
Table 3, along with Figure 7, shows the comparison. Both our regularization and weight subset optimization help unlearn the target image more effectively, without forgetting other concepts.
Target Images (to forget) Related Images (to retain) Other Images (to retain)
MSE ($\uparrow$) CLIP ($\downarrow$) MSE ($\downarrow$) CLIP ($\uparrow$) MSE ($\downarrow$) CLIP ($\uparrow$)
K=500 0.054±0.004 0.72±0.012 0.039±0.0003 0.86±0.001 0.041±0.0003 0.79±0.001
K=1000 0.058±0.004 0.68±0.014 0.041±0.0003 0.86±0.001 0.041±0.0003 0.79±0.001
K=4000 0.060±0.004 0.61±0.014 0.046±0.0004 0.83±0.001 0.041±0.0003 0.79±0.001
Table 4: Does leave-K-out models forget other images? We verify that leave-$K$-model forgets concepts specific to the target query. We report deviations (MSE, CLIP) from the target image, related images that are similar to the target, and other images of unrelated concepts. We find that target images deviate more than related and other images, while other images stay almost the same. Related images’s errors increase with larger $K$, but they are much smaller than target images’ deviations.
Figure 8: Does leave-$K$-out models forget other images? We show that leave-$K$-out model forgets the specific target (left), while retaining its generation on related images (middle) and images of other concepts (right).
Does leave-K-out models forget other images?   In Sec. 5, we show that leave-$K$-out models forget how to generate target synthesized image queries. Here we ask whether these models forget unrelated images, too. We study how much leave-$K$-out model’s generation deviates from those of the original model in three categories:
  • Target images: the attributed synthesized image. Leave-$K$-out models should forget these—the greater the deviation, the better.
  • Related images: images synthesized by captions similar to the target prompt. We obtain the most similar 100 captions from the MSCOCO val set using CLIP’s text encoder. Leave-$K$-out models should not forget all of them—the smaller the deviation, the better.
  • Other images: images of unrelated concepts. Prompts are 99 different captions selected from the MSCOCO val set. Leave-$K$-out models should not forget these—the smaller the deviation, the better.
In Fig 8, we find that the leave-$K$-out model “forgets” the query bus image specifically while retaining other buses and other concepts. In Table 4, we quantitatively report deviations using mean square error (MSE) and CLIP similarity, where we evaluate 40 pairs of target images and leave-K-out models. We observe that target images have larger MSE and lower CLIP similarity than related images and other images. Also, as the number of removed influential images ($K$) increases, the target image error increases rapidly while other images stay almost the same. Interestingly, related images’ errors increase with larger $K$, but the errors are still much smaller than those of target images. As $K$ increases, the group of influential images can start affecting other related concepts.
[…]
Method Loss K=500 K=1000 K=4000 K=500 K=1000 K=4000 K=500 K=1000 K=4000
TRAK $||\epsilon_{\theta}(\mathbf{x}_{t},t)-\epsilon||^{2}$ 5.2±0.14 5.8±0.16 7.1±0.16 4.7±0.21 4.7±0.24 4.7±0.20 7.6±0.09 7.6±0.09 7.5±0.09
D-TRAK $||\epsilon_{\theta}(\mathbf{x}_{t},t)||^{2}$ 5.4±0.16 6.6±0.22 9.6±0.33 5.9±0.24 6.4±0.25 7.8±0.30 7.3±0.10 7.1±0.09 6.4±0.09
[…]
Ours (D-TRAK) $||\epsilon_{\theta}(\mathbf{x}_{t},t)||^{2}$ 4.4±0.14 5.1±0.20 7.6±0.36 4.7±0.23 4.9±0.24 6.0±0.28 7.6±0.09 7.4±0.09 6.3±0.09
Table 5: Choice of loss function. We study different choices of loss. While D-TRAK [14] empirically showed that changing the diffusion loss to a square loss increases performance drastically for influence function, we find that this loss change does not grant performance gain for our formulation. We report $\Delta\mathcal{L}(\hat{\mathbf{z}},\theta)$ and $\Delta G_{\theta}(\epsilon,\mathbf{c})$ as in Table 1.
Applying the same loss function changes as D-TRAK.   As mentioned in Sec. 5.1, D-TRAK yields a significant performance gain by changing the loss function in TRAK from a denoising loss ($\mathbb{E}[\|\epsilon_{\theta}(\mathbf{x}_{t},\mathbf{c},t)-\epsilon\|^{2}]$) to a square loss ($\mathbb{E}[\|\epsilon_{\theta}(\mathbf{x}_{t},\mathbf{c},t)\|^{2}]$). In Table 5, we find that applying the same loss change to our method leads to worse performance.
@@ C.2 Customized Model Benchmark
Appendix D Efficiency Comparison with Other Methods
We compare our method’s efficiency with other methods. Below, we briefly overview each method:
  • Our unlearning method: we obtain a model that unlearns the synthesized image query and checks the loss increases for each training image after unlearning.
  • TRAK, JourneyTRAK: precompute randomly-projected loss gradients for each training image. Given a synthesized image query, the authors first obtain its random-projected loss gradient and then match it with the training gradients using the influence function. For good performance, both methods require running influence function on multiple pre-trained models (e.g., 20 for MSCOCO).
  • D-TRAK: The runtime complexity is similar to that of TRAK, but it is running the influence function on a single pre-trained model.
  • Feature similarity: standard image retrieval pipeline. Precompute features from the training set. Given a synthesized image query, we compute feature similarity for the entire database.
We compare the efficiency of each method.
Feature similarity is the most run-time efficient   since obtaining features is faster than obtaining losses or gradients from a generative model. However, feature similarity does not leverage knowledge of the generative model. Our method outperforms various feature similarity methods (Table 1).
Our method is more efficient than TRAK/JourneyTRAK/D-TRAK in precomputation.   Our method’s precomputation is much more efficient runtime and storage-wise compared to TRAK, JourneyTRAK, and D-TRAK. Our method only requires computing and storing loss values of training images from a single model. On the other hand, TRAK-based methods require pre-training extra models (e.g., 20) from scratch, and precomputing/storing loss gradients of training images from those models.
Our method is less efficient when estimating attribution score.   TRAK-based methods obtain attribution scores by taking a dot product between synthesized image gradient and the stored training gradient features. The methods require calculating and averaging such dot product scores from the 20 pretrained models. On the other hand, as acknowledged in our limitations (Sec. 6), though our model unlearns efficiently (e.g., only 1-step update for MSCOCO), getting our attribution score involves estimating the loss on the training set, which is less efficient than dot product search. Tradeoff-wise, our method has a low storage requirement at the cost of higher runtime.
Our main objective is to push the envelope on the difficult challenges of attribution performance. Improving computation efficiency of attribution is a challenge shared across the community [16, 29], and we leave it to future work.
@@ Appendix D Efficiency Comparison with Other Methods
Appendix E Change log.
v1   Initial release.
v2   NeurIPS 2024 camera ready. Figure 1 is updated with more training images. Section 5, Figures 235610111213, and 14 are edited to incorporate the two additional baselines. Tables 1 and 2 are added to clarify the experimental results. Additional analyses in the unlearning and leave-K-out models are added in Appendix C.1, Figures 78, and Tables 34. Additional analysis of D-TRAK is added in Appendix C.1, and efficiency comparison is added in Appendix D. Errors corrected in Appendix A.3.